OpenStudy (anonymous):

Find the missing leg of the right triangle if 7 is the other leg and 25 is the hypotenuse. 18 25.96 16 24

OpenStudy (anonymous):

You may get a more accurate answer if you post this to the mathematics subject!

OpenStudy (nurali):

Pythagorean Theroem: (7)^2 + (b)^2 = (25)^2 49 + b^2 = 625 b^2 = 576 b = 24